

********************************************************************
UARS WEEKLY STATUS REPORT - 21 JAN 2000
********************************************************************
***********
SOLSTICE
***********
We continue to gather enough data to meet all the SOLSTICE science
objectives during this period of reduced tape recorder operations.
SOLSTICE operates normally, collecting solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible.
***********
SUSIM
***********
SUSIM continues to gather a minimum set of daily solar scans scheduled
for periods of predicted telemetry.
Given the present situation of intermittent and uncertain UARS telemetry,
we are working to generate a version of the SUSIM production software
which is more robust with respect to incomplete scanning data. These
fixes are being tested now.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM/AXIS was activated on January 15 in the half AXIS mode. Only
AXIS-1 is active, producing energy deposition from electrons inferred from the
measured X-rays from eight pixels on one side of the spacecraft.
The PEM spectrometers continue to be active under PMON control*. While
under PMON control, the PEM spectrometers computed energy deposition from
electrons and protons measured in-situ by the particle spectrometers.
PEM submitted its paper discussing its arguments for the extended UARS
mission which would overlap with the CHEM spacecraft. PEM arguments for
extending the UARS mission included the possibility of coincidence measurements
of in situ particle measurements with remote-instruments measurements from CHEM,
a continuation of particle monitoring into the next solar cycle, and increasing
the time base where large storms are more likely to occur and dump energy into
the atmosphere.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
The MPG participated in a meeting on 20 January which was
called to make recommendations for spacecraft tape recorder
(NBTR2) scheduling. While several options are being explored,
NBTR2 will be used as needed to support specific science
mission or engineering requirements (using stored commands).
Real time data will continue to be relayed via TDRS E and
TDRS W when either TDRS is in view.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
31 days WIN (5.11A) PROD (L1-L3) (520R-311C-13F=196L)
2 days SEN (5.0) PROD (1992) (2R-2C=0L)
1 day SUS (7.2) PROD (L1-L3) (TEST) (10R-10C=0L)
27 days SUS (7.2) PROD (L1-L3) (27R-27C=0L)
The CDHF is processing SENSOR V5.0 software against 1992 data as requested
by FDF.
The CDHF has promoted SUSIM V7.2 Formal Software Delivery (Patch) to
production mode. Current day processing has resumed.
CDHF REHOST:
- Testing HRDI V4.4 on the Rehost System.
- Testing SOLSTICE V4.62 on the Rehost System.
- Plans to test WINDII V5.11A on the Rehost System.
- Rehost procedures/work instructions continued to be developed and/or
reviewed.
***********
FOT
***********
REPORTING PERIOD
This report covers 08 January, 2000 (Orbit 45499, GMT Day 008) through 14
January 2000 (Orbit 45603, GMT Day 014).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 009/1337 45522 Calibration
HALOE 009/1759 45525 On
The beta angle ranged between the angles of 2.2 degrees to 30.7 degrees for
this report period. The beta angle is now increasing toward a maximum angle of
76.0 degrees on 27 January 1999 (DOY 027).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.0 to 1.3 degrees C. Additional battery performance data for each
battery on 07 December 2000 and 14 December 2000 are provided in Appendix D.
The clock error ranged from -7.9 to 8.2 msec during this report period. Clock
rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% OUT-OF-LIMITS = 01
% ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT = 08
% HALOE = 02
% PEM = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 00
Three (3) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
7 Transponder A: 99/358 - 2287.496855
7 Transponder B: 99/365 - 2287.499111
GROUND SYSTEM OPERATION
One (1) operational support problems occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/013 show the observatory data loss to be 563 hours, 06
minutes, 50.601 seconds (an increase of 01 hours 10 minutes 14.240 seconds
since the last report period).
The increase of 01 hours 10 minutes 14.240 seconds occurred over six (6) days:
- 03 minutes 11.696 seconds on Day 007 due to an unspecified problem.
- 02 minutes 47.408 seconds on Day 009 due to an unspecified problem.
- 21 minutes 44.032 seconds on Day 010 due to an unspecified problem.
- 20 minutes 47.296 seconds on Day 011 due to an unspecified problem.
- 14 minutes 24.512 seconds on Day 012 due to an unspecified problem.
- 07 minutes 19.296 seconds on Day 013 due to an unspecified problem.
These calculations are under repair. They are being recalculated from the date
where PASS 2 operations were scheduled exclusively.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week we executed the PEM calibration. Everything went according to plan
and they were put back into day only operations after two orbits of full time
activity.
The FOT also continued their investigation of the TMONs so that the new TMON
can be built for the OBC controlled safehold. We got on the simulator and made
note of all the telemetry points that didn't match the real spacecraft. This
list will be given to UTTS manager so that they can be put into the scenario.
Once this is accomplished the writing and testing of the TMON will begin.
MEETINGS
JAN 12, 2000 - The FOT attended the UARS Monthly meeting. This was the first
meeting with our new Mission Director, Herman Williams. Also Mr Bill Guit was
present. Mr Guit is the Mission Director to Landsat 7 and TRMM as well as
fulfilling a backup roll to Herman. The FOT would like to welcome Mr. Williams
and Mr Guit to the project.
Aside from the usual subsystem overview most of the discussion were related to
the tape recorder. The data loss calculations are going to be recalculated
based on Pass 2 recordings only. The start date is to be determined.
Additional discussion centered on the components necessary for going to three
TDRSS operations. A meeting was scheduled later this month to discuss this
further.
Jan 11, 2000 - The FOT attended a meeting to brief the Flight Software Group
from Code 582. They will be assuming the responsibility for all spacecraft
software maintenance. If three TDRSS operations are approved, the FSW group
will probably be given that task of modifying the onboard code.
FUTURE SIGNIFICANT EVENTS
No specific date was agreed upon for UARS Safehold configuration discussions.
Instead, the Supporting Engineers involved will be contacted every two weeks to
determine if either the Flight Operations Team or Supporting Engineers have any
new developments that should be discussed in a group.
The next UARS Yaw slew is scheduled for February 15, 2000 (DOY 046). This
maneuver will be a Reverse to Forward Flight maneuver (Southern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 14 January 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
45522 009/1247 00-004 NBTR NBTR NOOPS
45522 009/1337 00-005 PEM PEM CALIBRATION
45525 009/1759 00-006 NBTR NBTR NOOPS
45525 009/1759 00-007 HALOE Power ON
45527 009/2127 00-008 HALOE Power ON
45541 010/1837 00-009 NBTR NBTR NOOPS
45555 011/1808 00-010 NBTR NBTR NOOPS
45564 012/0802 00-010 NBTR NBTR NOOPS
45570 012/1822 00-011 NBTR NBTR NOOPS
45585 013/1815 00-012 NBTR NBTR NOOPS
45600 014/1818 00-013 NBTR NBTR NOOPS
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 14 January 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPTSPBIMON Y-HI HRDI PB CURRENT MONITOR continuous
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
No Alerts to report this report period.
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 14 January 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
None this report period
_______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 14 January 2000
07 Jan 2000 - GMT Day 007 Beta = 2.2 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- --------- ----------- ----- ------- ------
1 44.8/-72.8 -3.68/-4.62 +0.0/+0.0 16.2 0.000 0.0
2 +16.8/-28.0 +5.79/+4.51 30.4/-13.6 24.5 1.030 21.4
3 +33.6/-22.4 +4.51/+3.23 29.6/-14.4 24.5 1.027 21.2
14 Jan 2000 - GMT Day 014 Beta = 30.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------- ---------- --------- ----- ------ ------
1 -50.4/-78.4 -3.05/-4.31 +0.0/+0.0 16.2 0.000 0.0
2 +11.2/-78.4 +6.44/+4.83 +30.4/-15.2 24.2 1.023 22.5
3 +33.6/-67.2 +5.15/+3.87 +29.6/-16.0 24.2 1.017 22.3
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
no changes V/T 5
__________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 14 January 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
No new reports generated
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
No new reports generated
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
No new reports generated
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-064 FOT 44135 10/8-9/99 N/A HRDI Cal Lamp Errors
99-071 FOT 44398 299/1212 21512 TPF Not Receiving 512 PB
00-001 FOT See list January 14170 Generic Late Acquisition
SEVNTFW00/02